Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by gmatassistance » Sun Mar 28, 2010 4:57 pm
A machine consists of three components, which are components X, Y, and Z. These three components operate properly independently of one another. The machine operates properly only when component X operates properly and at least one of components Y and Z operates properly. The probability that component X operates properly is 0.8, the probability that component Y operates properly is 0.4, and the probability that component Z operates properly is 0.3. What is the probability that the machine operates properly?

0.096
0.252
0.464
0.583
0.648
Join the discussion
Source: — Problem Solving |

by gmatassistance » Sun Mar 28, 2010 4:59 pm
I tried using the probability formula where the Probability of A or B = pA x pB - (pA + pB) to figure out the probability for components Y or Z to operate properly however got a negative answer.

Can someone please explain why the above formula didn't work and what other strategies to keep in mind for such questions?

Thanks!
Join the discussion

by student22 » Sun Mar 28, 2010 7:57 pm
I'm not sure if this is the right way to solve this, but couldn't you take the probability of X occurring, 0.80 * (1 - PZ and PY NOT OCCURRING). 0.80 * (1 - (.7 * .6) = 0.80 * .58 = .464.

My reasoning is that its easier to find the the probability that they will both fail and then subtract that from 1. This gives the probability of 1 work or both working, which you multiply by X.

Hopefully this helps, if not, maybe someone has a better way.
Join the discussion

by pops » Sun Mar 28, 2010 9:53 pm
machine will work properly if
(A) works and (B or C) works
this means,
p = A On, B On, C On + A On, B On, C Off + A On, B Off, C On
= 0.8*0.4*0.3 + 0.8*0.4*(1-0.3) + 0.8*(1-0.4)*0.3
= 0.464
Join the discussion

by koolk86 » Sun Mar 28, 2010 10:13 pm
I have solved it like this :

For the machine to work properly X must work properly and ATLEAST Y or Z must work properly

Probability of X working properly = 0.8

Probability of Y working properly = 0.4

Probability of z working properly = 0.3

Probability of X not working properly = 0.2 (1- 0.8)

Probability of Y not working properly = 0.6 (1- 0.4)

Probability of Z not working properly = 0.7 (1- 0.3)


Now use the above:

P(X works) * P(Y works) * P (Z does not work) = 0.8 * 0.4 * 0.7 = 0.244

OR

P(X works)* P(Ydoes not work) * P(Z works) = 0.8 *0.6*0.3 = 0.144

OR

P(X works) * P ( Y works ) * P ( Z works) = 0.8 * 0.4 * 0.3 = 0.096

Add All you get : 0.464

This is a bit long method and involves calculations.

I guess one could also solve by 1-x method ???
Join the discussion

by hooliganpete » Mon Mar 29, 2010 12:56 pm
I found this method to be quick and painless:

Machine works when X works and either Y or Z so find the probability that Either Y or Z works. The 1- method is great for this first step.

Prob(X doesn't work) * Prob(Y doesn't work) = .42

1-.42 = .58

Thus, .58 = Prob(either X or Y works)

Now since X has to work, .42*.8 = Prob(X works)*Prob(Either X or Y work) = .464
Join the discussion

by eaakbari » Mon Mar 29, 2010 12:58 pm
Hey gmatassistance
I think the mistake you are making here is applying the probability law.
It states that all the events are independent.
The law which you are applying is only applicable for non - independent events, even mutually exclusive events would do but nope not independent.

For independent remember the multiplication rule which is

For A and B to both occur that equals P(A) *P(B)

For this question realise that there are 3 cases where machine will work
X and Y are operating but Z is not
X and Z are operating but Y is not
X,Y and Z are operating

So prob will be
0.8*0.4*0.7 + 0.8*0.3*0.6 + 0.8*0.3*0.6

Which equals 0.464

Hope this helps

E
Join the discussion